OpenStudy (anonymous):

Evalutate the expression x^-2y^2/x^4y^-3 for x=4 and y=-2 a.-16 b.-1/16 c.-128 d.-1/128

OpenStudy (superdavesuper):

Put x=4 and y=-2 into the eqn....

OpenStudy (anonymous):

i tried but i don't understand can u help me @superdavesuper

OpenStudy (superdavesuper):

x=4 so what is x^-2 = ?

OpenStudy (anonymous):

idk can u just tell me please @superdavesuper

OpenStudy (superdavesuper):

i can tell u but it wont do u any good as u will not know how to do a similar question

OpenStudy (anonymous):

i think it is -2? @superdavesuper

OpenStudy (anonymous):

jk no its -16 right

OpenStudy (superdavesuper):

close but no... x=4; then x^-2 = 4^-2 = 1/4^2 = 1/16

OpenStudy (anonymous):

oh okay thank u @superdavesuper

OpenStudy (superdavesuper):

welcome n thanx 4 the medal! u can do the rest right?